OpenStudy (anonymous):

A certain voltage can be expressed as v(t)=(5+t^2)u(t-1). What is the value of this voltage at t=3sec?

OpenStudy (anonymous):

v(t)=(5+t^2)u(t-1) and you are give t =3, just pulg in 3 where evr u see the't' v(3)=(5+3^2)u(3-1) = (5+9)u(2) =28u

OpenStudy (shadowfiend):

Remember to give people a medal when they're helpful :)